How do you work this out? I feel real stupid but it has been a long time for me.
n(n-1)/2 = 21

Answer by stanbon(75887)

n(n-1)/2 = 21
-----
Multiply both sides by 2 to get:
n(n-1) = 42
---
Expand the left side
n^2 - n = 42
-----
Subtract 42 from both sides:
n^2 - n - 42 = 0
-----
Factor:
(n-7)(n+6) = 0
----
n-7 = 0 or n+6 = 6
n = 7 or n = -6
=====================
Cheers,
Stan H.
===========================